# Comprehensive Cognitive LLM Agent for Smartphone GUI Automation

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable potential as human-like autonomous language agents to interact with real-world environments, especially for graphical user interface (GUI) automation. However, those GUI agents require comprehensive cognition ability including exhaustive perception and reliable action response. We propose Comprehensive Cognitive LLM Agent, CoCo-Agent, with two novel approaches, comprehensive environment perception (CEP) and conditional action prediction (CAP), to systematically improve the GUI automation performance. First, CEP facilitates the GUI perception through different aspects and granularity, including screenshots and complementary detailed layouts for the visual channel and historical actions for the textual channel. Second, CAP decomposes the action prediction into sub-problems: action type prediction and action target conditioned on the action type. With our technical design, our agent achieves new state-of-the-art performance on AITW and META-GUI benchmarks, showing promising abilities in realistic scenarios. Code is available at https://github.com/xbmxb/AAgent.
Comprehensive Cognitive LLM Agent for Smartphone GUI Automation
Xinbei Ma, Zhuosheng Zhang, Hai Zhao School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University Department of Computer Science and Engineering, Shanghai Jiao Tong University Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University sjtumaxb@sjtu.edu.cn, zhangzs@sjtu.edu.cn, zhaohai@cs.sjtu.edu.cn
@@ 1 Introduction
Graphical user interface (GUI) automation aims to enable human-like operations on operating systems with artificial intelligence instead of human efforts. Large language models (LLMs) have demonstrated commendable performance as human-like agents, showing emergent abilities of perceiving Yao et al. (2022), reasoning Li et al. (2023b); Park et al. (2023), and acting Wang et al. (2023); Richards (2023). With multi-modal enhancement, LLM agents become promising autonomous GUI assistants to deal with complex tasks on behalf of human operators. To interact with the GUI environment, those agents require comprehensive cognition, including exhaustive perception and reliable action response.
Current vital challenges for autonomous agents lie in two aspects. One is (i) the dependence on strong LLMs, and the other is (ii) the insufficient GUI environment modeling.
Although strong LLMs like GPT-4V OpenAI (2023) and ChatGPT Ouyang et al. (2022) ignite the development of autonomous agents, they exhibit shortcomings in realistic use. First, the alignment requires a trustworthy domain transfer as there is a large disparity between GUI commands and natural languages. GUI agents are expected to generate accurate and well-formulated responses as executable GUI commands, which is non-trivial for zero-shot prompting. For example, given the GUI that accepts commands as “{action: click, touch_point:[y, x], touch_point:[y, x], typed_text: ‘’}”, semantically equivalent generations like “Open the address book on your phone” is plausible but unavailable. Second, the black-box APIs are likely to cause unexpected safety issues. Risks to privacy and integrity may arise when granting personal device authority to a black-box API. This significantly reduces realistic usability. Third, the performance mainly relies on the prompt design. The issues mentioned above leave heavy burdens on the design of prompt lines for those agents. Besides necessary environment descriptions, the prompts (and post-processing) need to be sophisticated to enhance domain alignment, instruction following, and security risk mitigation in different circumstances.
@@
This work proposes CoCo-Agent, a Comprehensive Cognitive LLM Agent, to address the challenges above for smartphone GUI automation. CoCo-Agent adopts a multimodal backbone of LLaVA Liu et al. (2023) and further enhances comprehensive cognition, respectively for exhaustive perception and reliable action response. The two proposed approaches are comprehensive environment perception (CEP) and conditional action prediction (CAP). Specifically, CEP integrates GUI perception elements of textual goal, historical action, and high-level and detailed description of the vision channel. CAP decomposes the complex and redundant GUI action commands into sub-problems following a top-down order. Our experiments cover diverse tasks in two GUI benchmarks, AITW Rawles et al. (2023) and META-GUI Sun et al. (2022a), including application manipulation, web operation, and dialogues. CoCo-Agent achieves SOTA performance with a limited parameter size. Subsequently, we present deep analyses including element ablation, visual module selection, and future action prediction. We show the significant effect of each perception element and the favorable choice of visual module. We also analyze the limitations of existing datasets and illustrate the additional potential of CoCo-Agent for realistic scenarios.
@@ 2.1 Autonomous Language Agents
One research line relies on the strong fundamental competencies of LLMs. Based on ChatGPT or GPT-4, autonomous agents can be built by only well-written prompts. Existing works have proved the reasoning, planning, and generalizing abilities of GPT-based agents, e.g., AutoGPT Richards (2023), BabyAGI Nakajima (2023), AgentGPT Reworkd (2023), HuggingGPT Shen et al. (2023), and MM-Navigator Yan et al. (2023).
@@ 3.1 Task Formalization
PRESS_HOME |
"[-1.0, -1.0]" |
"[-1.0, -1.0]" |
I need to <PRESS_HOME> |
PRESS_BACK |
"[-1.0, -1.0]" |
"[-1.0, -1.0]" |
I need to <PRESS_BACK> |
PRESS_ENTER |
"[-1.0, -1.0]" |
"[-1.0, -1.0]" |
I need to <PRESS_ENTER> |
STATUS_TASK_COMPLETE |
"[-1.0, -1.0]" |
"[-1.0, -1.0]" |
For this goal, no more action is needed, so <STATUS_TASK_COMPLETE> |
TYPE |
"[-1.0, -1.0]" |
"[-1.0, -1.0]" |
"{string}" |
I need to <TYPE> a string here, "typed_text": "{string}" |
DUAL_POINT |
"{coordinate}" |
"{coordinate}" |
"{string}" |
I need to <SCROLL> {direction} |
DUAL_POINT |
"{coordinate}" |
"{coordinate}" |
"{string}" |
I need to <CLICK> {item name}, the location of {item name} is "tap_point": "{coordinate}" |
DUAL_POINT |
"{coordinate}" |
"{coordinate}" |
"{string}" |
I need to <TAP> on the screen, the location is "tap_point": "{coordinate}" |
@@ 4.1 Dataset
General |
9,476 |
85,413 |
545 |
Install |
25,760 |
250,058 |
688 |
Google Apps |
625,542 |
4,903,601 |
306 |
Single |
26,303 |
85,668 |
15,366 |
Web Shopping |
28,061 |
365,253 |
13,473 |
@@ 4.2 Implementation
AITW |
Overall | General | Install | GoogleApps | Single | WebShop. |
PaLM-2 |
39.6 |
ChatGPT |
7.72 |
5.93 |
4.38 |
10.47 |
9.39 |
8.42 |
MM-Navigator |
50.54 |
41.66 |
42.64 |
49.82 |
72.83 |
45.73 |
MM-Navigator w/ text |
51.92 |
42.44 |
49.18 |
48.26 |
76.34 |
43.35 |
MM-Navigator w/ history |
52.96 |
43.01 |
46.14 |
49.18 |
78.29 |
48.18 |
@@
68.7 |
BC w/ history |
73.1 |
63.7 |
77.5 |
75.7 |
80.3 |
68.5 |
LLaMA-2 |
28.40 |
28.56 |
35.18 |
30.99 |
27.35 |
19.92 |
Auto-UI |
74.22 |
65.94 |
77.62 |
76.45 |
81.39 |
69.72 |
Auto-UI |
74.27 |
68.24 |
76.89 |
71.37 |
84.58 |
70.26 |
CogAgent |
76.88 |
65.38 |
78.86 |
74.95 |
93.49 |
71.73 |
@@
70.37 |
58.93 |
72.41 |
70.81 |
83.73 |
65.98 |
CoCo-Agent |
76.96 |
65.60 |
80.60 |
75.76 |
88.81 |
74.02 |
CoCo-Agent |
79.05 | 70.96 | 81.46 | 76.45 |
91.41 |
75.00 |
@@
General |
LLaVA |
58.93 |
80.08 |
N/A |
56.76 |
63.31 |
93.29 |
CoCo-Agent |
70.96 |
87.49 |
76.72 |
68.91 |
75.80 |
97.10 |
@@
LLaVA |
65.98 |
85.43 |
N/A |
64.81 |
68.61 |
92.60 |
CoCo-Agent |
76.10 |
89.80 |
80.20 |
73.88 |
78.48 |
96.96 |
META-GUI |
Action | Act. type | Item | Direction | Input (F1) | Utter. (BLEU) |
LayoutLM |
67.76 |
82.22 |
71.98 |
94.87 |
90.56 |
50.43 |
LayoutLMv2 |
64.48 |
85.60 |
64.38 |
92.95 |
70.76 |
58.20 |
BERT |
78.42 |
87.52 |
82.84 |
93.59 |
97.24 |
62.19 |
m-BASH |
82.74 |
90.80 |
85.90 |
96.42 |
94.23 |
63.11 |
@@
76.27 |
87.47 |
77.49 |
98.18 |
96.06 |
67.24 |
LLaVA w/ history |
81.08 |
91.68 |
81.23 |
97.62 |
96.93 |
66.57 |
CoCo-Agent |
88.27 | 92.59 | 91.72 | 98.39 |
96.15 |
65.90 |
@@ 4.3 Baselines
Multi-modal API-based methods. MM-Navigator Yan et al. (2023) is a GPT-4V-based multimodal agent, achieving few-shot SOTA.
@@ 5.1.1 Ablation
Our method combines CEP and CAP to characterize the GUI environment. We ablate each CEP element along with the refactoring method of CAP to observe the significance for the CoCo-Agent. Results are shown in Table 5.222The first 1000 samples in General, the most difficult subset of AITW, are tested for our extensive analysis. The improvement of each element is significant, especially for layouts (+7.7%) and action history (+5.1%).
@@
Goal |
Img |
CAP |
Layout |
History |
General | Meta |
0 |
58.40 |
73.60 |
0 |
61.00 |
76.90 |
0 |
68.70 |
85.55 |
8 types |
71.00 |
86.99 |
8 actions |
73.80 |
88.27 |
@@ 5.1.2 Replacement
Replace |
None |
Goal |
Img |
Layout |
Hist. |
Avg. |
Detection |
N/A |
94.60 |
93.94 |
91.02 |
92.69 |
93.06 |
Action |
70.96 |
63.21 |
59.69 |
57.55 |
58.45 |
59.73 |
@@
Model |
Visio Encoder |
LM |
Acc |
Auto-UI |
BILP-2 Encoder |
FLAN-Alpaca |
65.9 |
MMICL2 |
Q-Former |
Flan-T5-xxl |
56.4 |
mPLUG2 |
Abstractor |
Vicuna-7B |
53.0 |
CogAgent |
Low & High Res. |
Vicuna-7B |
65.4 |
LLaVA |
CLIP |
Llama-2-chat-7B |
58.9 |
Ours |
CLIP & Layout |
Llama-2-chat-7B |
71.0 |
Train |
1-next |
3-next |
3-next |
3-next |
Test |
1-next |
1-next |
2-next |
3-next |
Accuracy |
70.96 |
58.90 |
43.40 |
37.74 |
@@
Prop. |
Acc |
Prop. |
Acc |
Prop. |
Acc |
Prop. |
Acc |
Prop. |
Acc |
Prop. |
Acc |
General |
LLaVA |
86.09 |
54.16|84.71 |
10.90 |
86.09 |
1.17 |
8.16 |
5.36 |
77.73 |
2.61 |
41.55 |
10.67 |
62.53 |
CoCo-Agent |
67.73|91.59 |
84.34 |
15.31 |
89.76 |
57.99 |
78.19 |
@@ 5.2 Visual Capability
The results are shown in Table 7. Our integrated CLIP with projector encodes an image to a 256-length vector with a 4096 hidden size.1 With fine-grained layouts, CoCo-Agent receives exhaustive visual information. Auto-UI uses a BLIP-2 with pooling leading to a 1-length image vector, which is integrated into language embedding by an attention-based fusion module. Differently, MMICL adopts Q-Former to learn a 32-length query vector, and mPLUG adopts Abstractor to learn a 64-length vector. CogAgent uses an EVA2-CLIP-E Sun et al. (2023) as a low-resolution encoder and an EVA2-CLIP-L Sun et al. (2023) with cross-attention layers as a high-resolution encoder. In cross-attention layers, the high-resolution image features interact with each layer of the language model.

### v3

# CoCo-Agent: A Comprehensive Cognitive MLLM Agent for Smartphone GUI Automation

@@ Abstract
Multimodal large language models (MLLMs) have shown remarkable potential as human-like autonomous language agents to interact with real-world environments, especially for graphical user interface (GUI) automation. However, those GUI agents require comprehensive cognition ability including exhaustive perception and reliable action response. We propose a Comprehensive Cognitive LLM Agent, CoCo-Agent, with two novel approaches, comprehensive environment perception (CEP) and conditional action prediction (CAP), to systematically improve the GUI automation performance. First, CEP facilitates the GUI perception through different aspects and granularity, including screenshots and complementary detailed layouts for the visual channel and historical actions for the textual channel. Second, CAP decomposes the action prediction into sub-problems: action type prediction and action target conditioned on the action type. With our technical design, our agent achieves new state-of-the-art performance on AITW and META-GUI benchmarks, showing promising abilities in realistic scenarios. Code is available at https://github.com/xbmxb/CoCo-Agent.
CoCo-Agent: A Comprehensive Cognitive MLLM Agent
for Smartphone GUI Automation
Xinbei Ma1,2,3,4, Zhuosheng Zhang1,∗, Hai Zhao1,2,3,4††thanks: Corresponding authors. This research was supported by the Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400), the Joint Funds of the National Natural Science Foundation of China (Grant No. U21B2020). 1School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University 2Department of Computer Science and Engineering, Shanghai Jiao Tong University 3Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University 4Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3 sjtumaxb@sjtu.edu.cn, zhangzs@sjtu.edu.cn, zhaohai@cs.sjtu.edu.cn
@@ 1 Introduction
Graphical user interface (GUI) automation aims to enable human-like operations on operating systems with artificial intelligence instead of human efforts. Large language models (LLMs) have demonstrated commendable performance as human-like agents, showing emergent abilities of perceiving Yao et al. (2022), reasoning Li et al. (2023b); Park et al. (2023), and acting Wang et al. (2023); Richards (2023). With the multimodal enhancement, MLLM agents become promising autonomous GUI assistants to deal with complex tasks on behalf of human operators. To interact with the GUI environment, those agents require comprehensive cognition, including exhaustive perception and reliable action response.
Current vital challenges for autonomous agents lie in two aspects. One is (i) the dependence on strong (M)LLMs, and the other is (ii) the insufficient GUI environment modeling.
Although strong (M)LLMs like GPT-4V OpenAI (2023) and ChatGPT Ouyang et al. (2022) ignite the development of autonomous agents, they exhibit shortcomings in realistic use. First, the alignment requires a trustworthy domain transfer as there is a large disparity between GUI commands and natural languages. GUI agents are expected to generate accurate and well-formulated responses as executable GUI commands, which is non-trivial for zero-shot prompting. For example, given the GUI that accepts commands as “{action: click, touch_point:[y0, x0], touch_point:[y1, x1], typed_text: ‘’}”, semantically equivalent generations like “Open the address book on your phone” is plausible but unavailable. Second, the black-box APIs are likely to cause unexpected safety issues. Risks to privacy and integrity may arise when granting personal device authority to a black-box API. This significantly reduces realistic usability. Third, the performance mainly relies on the prompt design. The issues mentioned above leave heavy burdens on the design of prompt lines for those agents. Besides necessary environment descriptions, the prompts (and post-processing) need to be sophisticated to enhance domain alignment, instruction following, and security risk mitigation in different circumstances.
@@
This work proposes CoCo-Agent, a Comprehensive Cognitive MLLM Agent, to address the challenges above for smartphone GUI automation. CoCo-Agent adopts a multimodal backbone of LLaVA Liu et al. (2023) and further enhances comprehensive cognition, respectively for exhaustive perception and reliable action response. The two proposed approaches are comprehensive environment perception (CEP) and conditional action prediction (CAP). Specifically, CEP integrates GUI perception elements of textual goal, historical action, and high-level and detailed description of the vision channel. CAP decomposes the complex and redundant GUI action commands into sub-problems following a top-down order. Our experiments cover diverse tasks in two GUI benchmarks, AITW Rawles et al. (2023) and META-GUI Sun et al. (2022a), including application manipulation, web operation, and dialogues. CoCo-Agent achieves SOTA performance with a limited parameter size. Subsequently, we present deep analyses including element ablation, visual module selection, and future action prediction. We show the significant effect of each perception element and the favorable choice of visual module. We also analyze the limitations of existing datasets and illustrate the additional potential of CoCo-Agent for realistic scenarios.
@@ 2.1 Autonomous Language Agents
One research line relies on the strong fundamental competencies of (M)LLMs. Based on ChatGPT or GPT-4, autonomous agents can be built by only well-written prompts. Existing works have proved the reasoning, planning, and generalizing abilities of GPT-based agents, e.g., AutoGPT Richards (2023), BabyAGI Nakajima (2023), AgentGPT Reworkd (2023), HuggingGPT Shen et al. (2023), and MM-Navigator Yan et al. (2023).
@@ 3.1 Task Formalization
PRESS_HOME | "[-1.0, -1.0]" | "[-1.0, -1.0]" | "" | I need to <PRESS_HOME> |
PRESS_BACK | "[-1.0, -1.0]" | "[-1.0, -1.0]" | "" | I need to <PRESS_BACK> |
PRESS_ENTER | "[-1.0, -1.0]" | "[-1.0, -1.0]" | "" | I need to <PRESS_ENTER> |
STATUS_TASK_COMPLETE | "[-1.0, -1.0]" | "[-1.0, -1.0]" | "" | For this goal, no more action is needed, so <STATUS_TASK_COMPLETE> |
TYPE | "[-1.0, -1.0]" | "[-1.0, -1.0]" | "{string}" | I need to <TYPE> a string here, "typed_text": "{string}" |
DUAL_POINT | "{coordinate}" | "{coordinate}" | "{string}" | I need to <SCROLL> {direction} |
DUAL_POINT | "{coordinate}" | "{coordinate}" | "{string}" | I need to <CLICK> {item name}, the location of {item name} is "tap_point": "{coordinate}" |
DUAL_POINT | "{coordinate}" | "{coordinate}" | "{string}" | I need to <TAP> on the screen, the location is "tap_point": "{coordinate}" |
@@ 4.1 Dataset
General | 9,476 | 85,413 | 545 |
Install | 25,760 | 250,058 | 688 |
Google Apps | 625,542 | 4,903,601 | 306 |
Single | 26,303 | 85,668 | 15,366 |
Web Shopping | 28,061 | 365,253 | 13,473 |
@@ 4.2 Implementation
AITW | API | Modality | Unified | Overall | General | Install | GoogleApps | Single | WebShop. |
PaLM-2 | PaLM-2 | Text | ✓ | 39.6 | – | – | – | – | – |
ChatGPT | ChatGPT | Text | ✓ | 7.72 | 5.93 | 4.38 | 10.47 | 9.39 | 8.42 |
MM-Navigator | GPT-4V | Text+Vision | ✓ | 50.54 | 41.66 | 42.64 | 49.82 | 72.83 | 45.73 |
MM-Navigator | GPT-4V | Text+Vision | ✓ | 51.92 | 42.44 | 49.18 | 48.26 | 76.34 | 43.35 |
MM-Navigator | GPT-4V | Text+Vision | ✓ | 52.96 | 43.01 | 46.14 | 49.18 | 78.29 | 48.18 |
@@
BC | N/A | Text+Vision | ✗ | 73.1 | 63.7 | 77.5 | 75.7 | 80.3 | 68.5 |
LLaMA-2 | N/A | Text | ✗ | 28.40 | 28.56 | 35.18 | 30.99 | 27.35 | 19.92 |
Auto-UI | N/A | Text+Vision | ✗ | 74.22 | 65.94 | 77.62 | 76.45 | 81.39 | 69.72 |
Auto-UI | N/A | Text+Vision | ✓ | 74.27 | 68.24 | 76.89 | 71.37 | 84.58 | 70.26 |
CogAgent | N/A | Text+Vision | ✗ | 76.88 | 65.38 | 78.86 | 74.95 | 93.49 | 71.73 |
@@
CoCo-Agent | N/A | Text+Vision | ✗ | 77.82 | 69.92 | 80.60 | 75.76 | 88.81 | 74.02 |
CoCo-Agent | N/A | Text+Vision | ✓ | 79.05 | 70.96 | 81.46 | 76.45 | 91.41 | 75.00 |
@@
General | LLaVA | 58.93 | 80.08 | N/A | 56.76 | 63.31 | 93.29 |
CoCo-Agent | 70.96 | 87.49 | 76.72 | 68.91 | 75.80 | 97.10 | |
@@
CoCo-Agent | 76.10 | 89.80 | 80.20 | 73.88 | 78.48 | 96.96 |
META-GUI | API | Modality | Action | Act. type | Item | Direction | Input (F1) | Utter. (BLEU) |
LayoutLM | N/A | Text | 67.76 | 82.22 | 71.98 | 94.87 | 90.56 | 50.43 |
LayoutLMv2 | N/A | Text+Vision | 64.48 | 85.60 | 64.38 | 92.95 | 70.76 | 58.20 |
BERT | N/A | Text | 78.42 | 87.52 | 82.84 | 93.59 | 97.24 | 62.19 |
m-BASH | N/A | Text+Vision | 82.74 | 90.80 | 85.90 | 96.42 | 94.23 | 63.11 |
@@
LLaVA | N/A | Text+Vision | 81.08 | 91.68 | 81.23 | 97.62 | 96.93 | 66.57 |
CoCo-Agent | N/A | Text+Vision | 88.27 | 92.59 | 91.72 | 98.39 | 96.15 | 65.90 |
@@ 4.3 Baselines
Multimodal API-based methods. MM-Navigator Yan et al. (2023) is a GPT-4V-based multimodal agent, achieving few-shot SOTA.
@@ 5.1.1 Ablation
Our method combines CEP and CAP to characterize the GUI environment. We ablate each CEP element along with the refactoring method of CAP to observe their significance for the CoCo-Agent. Results are shown in Table 5. The improvement of each element is significant, especially for layouts (+5.82%) and action history (+5.63%).
@@
Goal | Img. | CAP | Layout | History | General | META |
✓ | ✓ | ✗ | ✗ | 0 | 57.81 | 73.60 |
✓ | ✓ | ✓ | ✗ | 0 | 58.47 | 76.90 |
✓ | ✓ | ✓ | ✓ | 0 | 64.29 | 85.55 |
✓ | ✓ | ✓ | ✓ | 8 act. types | 67.80 | 86.99 |
✓ | ✓ | ✓ | ✓ | 8 full actions | 69.92 | 88.27 |
@@ 5.1.2 Replacement
Replace | None | Goal | Img | Layout | Hist. | Avg. |
Detection | N/A | 94.60 | 93.94 | 91.02 | 92.69 | 93.06 |
Action | 70.96 | 63.21 | 59.69 | 57.55 | 58.45 | 59.73 |
@@
Model | Vision Encoder | LM | Acc |
Auto-UI | BILP-2 Encoder | FLAN-Alpaca | 65.9 |
MMICL2 | Q-Former | Flan-T5-xxl | 56.4 |
mPLUG2 | Abstractor | Vicuna-7B | 53.0 |
CogAgent | Low & High Res. | Vicuna-7B | 65.4 |
LLaVA | CLIP | Llama-2-chat-7B | 58.9 |
Ours | CLIP & Layout | Llama-2-chat-7B | 71.0 |
Train | 1-next | 3-next | 3-next | 3-next |
Test | 1-next | 1-next | 2-next | 3-next |
Accuracy | 70.96 | 58.90 | 43.40 | 37.74 |
@@
Prop. | Acc | Prop. | Acc | Prop. | Acc | Prop. | Acc | Prop. | Acc | Prop. | Acc | ||
General | LLaVA | 86.09 | 54.16|84.71 | 10.90 | 86.09 | 1.17 | 8.16 | 5.36 | 77.73 | 2.61 | 41.55 | 10.67 | 62.53 |
CoCo-Agent | 67.73|91.59 | 84.34 | 15.31 | 89.76 | 57.99 | 78.19 | |||||||
@@ 5.2 Visual Capability
The results are shown in Table 7. Our integrated CLIP with projector encodes an image to a 256-length vector with a 4096 hidden size.1 With fine-grained layouts, CoCo-Agent receives exhaustive visual information. Auto-UI uses a BLIP-2 with pooling leading to a 1-length image vector, which is integrated into language embedding by an attention-based fusion module. Differently, MMICL adopts Q-Former to learn a 32-length query vector, and mPLUG adopts Abstractor to learn a 64-length vector. 222The first 1000 samples in General, the most difficult subset of AITW. CogAgent uses an EVA2-CLIP-E Sun et al. (2023) as a low-resolution encoder and an EVA2-CLIP-L Sun et al. (2023) with cross-attention layers as a high-resolution encoder. In cross-attention layers, the high-resolution image features interact with each layer of the language model.